How many molecules of ascorbic acid (vitamin c, C6 H8 O6) ARE in a 500mg tablet?
1
Expert's answer
2019-10-23T07:01:14-0400
"n(C_6H_8O_6) =\frac{m}{M} = \frac{0.5}{176.12} = 0.00284 mol"

"N = N_A\times n = 6.022\times 10^{23}\times 0.00284 =1.71\times 10^{21} molecules"